11. For the given reaction at a certain temperature, the equilibrium concentrations were found to be 2] 5.7x 10- M, [o2)- 1.o x from these data (delete units). O x 10 2M, and [NO] - 2.0 x 10 M. Calculate the value of the equilibrium constant and 2NO(g) + Olg)-2NO2(g) a. 8.1 x 102 b. 14.3 x 104 C. 2.9% 102 d. 1.2 x 103 e. None of these
Answer #1

For the given reaction, equilibrium constant

K = [NO2]^2/[NO]^2 *[O2] = (5.7*10^-3)^2/(2*10^-3)^2*(10^-2)

K = 32.49*10^-6 /4*10^-6*10^-2

K = 8.1225*10^2

Answer: a) 8.1*10^2